### v2

@@ 7 Conclusion
Motivated by a real-world use case, we initiated the study of parameterized local search heuristics for single-machine scheduling from a theoretical and practical point of view. Our experimental evaluations indicate that parameterized local search is a promising technique in the design of heuristics for the $1\mid\text{ST}_{\text{sd,f}}\mid C_{\max}$ scheduling problem. In comparison to the PILS1 algorithm [31], our experiments indicate that parameterized local search via windows of consecutive jobs can be a good alternative to random perturbation in a VNS. Besides the positive results, we also theoretically outlined the limits of parameterized local search via insert and swap operations.
There are many ways to extend our results in future work. Memetic algorithms [29] combine local search strategies with genetic algorithms. How do memetic algorithms work when the local search steps are performed with $k>1$? For which $k$ do they deliver a good trade-off between solution quality and running time? Considering the parameterized complexity, we may ask whether Swap LS MM and Insert LS MM are FPT for the sum of $k$ and the number $t$ of distinct types in the instance. Note that the number of types in the instance constructed in the proof of Theorem 4.1 is unbounded in the parameter. Thus, FPT for $k+t$ is still possible. This parameterization would be particularly interesting, since $t$ is often relatively small in industrial applications (recall that we have $t=8$ in our CUNA use case). Furthermore, it might be interesting to see whether the results in this work can be extended to other target functions than the makespan: is it possible to obtain similar results if one aims to minimize the (weighted) total tardiness or the maximum tardiness?

### v3

@@ 7 Conclusion
Motivated by a real-world use case, we initiated the study of parameterized local search heuristics for single-machine scheduling from a theoretical and practical point of view. Our experimental evaluations indicate that parameterized local search is a promising technique in the design of heuristics for the $1\mid\text{ST}_{\text{sd,f}},\overline{d_{j}}\mid C_{\max}$ scheduling problem. In comparison to the PILS1 algorithm [31], our experiments indicate that parameterized local search via windows of consecutive jobs can be a good alternative to random perturbation in a VNS. Besides the positive results, we also theoretically outlined the limits of parameterized local search via insert and swap operations.
There are many ways to extend our results in future work. Memetic algorithms [29] combine local search strategies with genetic algorithms. How do memetic algorithms work when the local search steps are performed with $k>1$? For which $k$ do they deliver a good trade-off between solution quality and running time? Considering the parameterized complexity, we may ask whether Swap LS MM and Insert LS MM are FPT for the sum of $k$ and the number $t$ of distinct types in the instance. Note that the number of types in the instance constructed in the proof of Theorem 5 is unbounded in the parameter. Thus, FPT for $k+t$ is still possible. This parameterization would be particularly interesting, since $t$ is often relatively small in industrial applications (recall that we have $t=8$ in our CUNA use case). Furthermore, it might be interesting to see whether the results in this work can be extended to other target functions than the makespan: is it possible to obtain similar results if one aims to minimize the (weighted) total tardiness or the maximum tardiness?